OpenStudy (anonymous):

Solve the following equation, giving the exact solutions which lie in [0, 2π). cos^3(x) = −cos(x)

OpenStudy (anonymous):

start with \[\cos^3(x)+\cos(x)=0\] then factor as \[\cos(x)(\cos^2(x)+1)=0\]

OpenStudy (anonymous):

since \(\cos^2(x)+1\) can never be zero, solve \(\cos(x)=0\) and you are done

OpenStudy (anonymous):

Should that be pi/2 and 3pi/2 ?

OpenStudy (anonymous):

why cos^2(x)+1 can never be zero anyway?

OpenStudy (anonymous):

@Rahaf, \(\cos^2x\ge0,\) so if you add one to both sides, you have \(\cos^2x+1\ge1\).

OpenStudy (anonymous):

oh ok thanks @SithsAndGiggles